OpenStudy (anonymous):

what is the simplified form of x^2+6x+8/4-x^2

OpenStudy (anonymous):

factor the numerator as \((x+2)(x+4)\) and the denominator as \((2-x)(2+x)\) then cancel the common factor

OpenStudy (anonymous):

3. x^2 + 6x + 8 / 4 - x^2? x+6x+2-x 6x+2 is your final answer.

OpenStudy (anonymous):

i am fairly sure the question is \[\frac{x^2+6x+8}{4-x^2}\]

OpenStudy (anonymous):

im confused. would yu do distributive property for the top?

OpenStudy (anonymous):

well yes, you would use the distributive property to factor. you get \[\frac{x^2+6x+8}{4-x^2}=\frac{(x+4)(x+2)}{(2-x)(2+x)}\] \[=\frac{(x+4)\cancel{(x+2)}}{(2-x)\cancel{(2+x)}}=\frac{x+4}{2-x}\]

OpenStudy (anonymous):

omg. okay thanks.

OpenStudy (anonymous):

yw